OpenStudy (anonymous):

Find all the zero's of f: f(t) = (t - 3)2 - (t + 3)

OpenStudy (tad1):

t = 3

OpenStudy (anonymous):

is it (t-3)2 or (t-3)^2 ?????

OpenStudy (tad1):

multiply it out and combine like terms

OpenStudy (tad1):

Sorry I had a sign error; the answer should be 9

OpenStudy (anonymous):

@tad I think the 2 in the first term shud b square.....

OpenStudy (anonymous):

if it is (t-3)^2 - (t+3), then = t^2 -6t +9 - t -3 (using the identity (a-b)^2 = a^2 - 2ab + b^2) = t^2 - 7t + 6 = t^2 - t - 6t + 6 = t(t-1) -6(t -1) = (t-6)(t-1)

OpenStudy (anonymous):

if f(t) = 0 => (t-6)(t-1) = 0 this is possible if either of the two factors is zero if t-6 = 0 => t = 6 if t-1 = 0 => t = 1 so zeros of f(t) are 1 and 6

OpenStudy (anonymous):

which one is correct nany?????????

OpenStudy (anonymous):

i got the las answer

OpenStudy (anonymous):

if u have to write square of something use the upward pointing carat ^ before the index eg 3 squared = 3^2 square of (x +9) = (x+9)^2 cube of (z-j) = (z-j)^3